OpenStudy (appleduardo):

what is the integral of e^(senx) 4cosx dx ? how can i solve it?

OpenStudy (appleduardo):

\[\int\limits_{}^{}e ^{sen x} 4\cos x dx\]

geerky42 (geerky42):

sen? You mean sec?

OpenStudy (anonymous):

try \(u=\sin(x), du=\cos(x)dx\) and you get it in one step

OpenStudy (appleduardo):

i got \[[e^{sen x} +c] [4 sen x + c]\] is that correct?

OpenStudy (appleduardo):

i meant "sin":

OpenStudy (tkhunny):

\(\int e^{\sin(x)}\cdot 4\cos(x)\;dx\) Following satellite73 suggestion u = sin(x) du = cos(x)dx This gives \(\int e^{u}\cdot 4\;du = 4\cdot e^{u} + C\) Substitute back to where we started. \(4\cdot e^{\sin(x)} + C\) Be careful, consistent, and confident.

OpenStudy (appleduardo):

thank you so much! but what happened with cos ?

OpenStudy (tkhunny):

It's all in there with the nature of the substitution. See the definition of du.

OpenStudy (appleduardo):

so in this case cos represents the derivative for sin in the formula , right?

OpenStudy (tkhunny):

That is where it came from. You can't just substitute a function. The nature of dx changes when you do that. Is English your first language? The answer to this question might help other folks understand where "sen(x)" came from.

OpenStudy (appleduardo):

haha yeah, uhmm but right now im studying in a spanish-speaking country, so sometimes (unconsciously) isay or write spanish :/ . thank you so!

OpenStudy (tkhunny):

No worries - as long as you don't mind freaking people out when you accidentally write the spanish versions of things. Good work!